### v2

# Safe and Generalized end-to-end Autonomous Driving System with Reinforcement Learning and Demonstrations

Abstract
An intelligent driving system should be capable of dynamically formulating appropriate driving strategies based on the current environment and vehicle status, while ensuring the security and reliability of the system. However, existing methods based on reinforcement learning and imitation learning suffer from low safety, poor generalization, and inefficient sampling. Additionally, they cannot accurately predict future driving trajectories, and the accurate prediction of future driving trajectories is a precondition for making optimal decisions. To solve these problems, in this paper, we introduce a Safe and Generalized end-to-end Autonomous Driving System (SGADS) for complex and various scenarios. Our SGADS incorporates variational inference with normalizing flows, enabling the intelligent vehicle to accurately predict future driving trajectories. Moreover, we propose the formulation of robust safety constraints. Furthermore, we combine reinforcement learning with demonstrations to augment search process of the agent. The experimental results demonstrate that our SGADS can significantly improve safety performance, exhibit strong generalization, and enhance the training efficiency of intelligent vehicles in complex urban scenarios compared to existing methods.
Figure 1: Overview of the safe and generalized end-to-end autonomous driving with reinforcement learning and demonstrations
1 Introduction
An intelligent autonomous driving systems must be able to handle complex road geometry and topology, complex multi-agent interactions with dense surrounding dynamic objects, and accurately follow the planning and obstacle avoidance. Current, autonomous driving systems in industry are mainly using a highly modularized hand-engineered approach, for example, perception, localization, behavior prediction, decision making and motion control, etc. (Thrun et al., 2006) and (Urmson et al., 2008). Particularly, the autonomous driving decision making systems are focusing on the non-learning model-based methods, which often requires to manually design a driving policy(González et al., 2015) and (Paden et al., 2016). However, the manually designed policy could have two several weaknesses: 1) Accuracy. The driving policy of human heuristics and pre-training model can be suboptimal, which will lead to either conservative or aggressive driving policies. 2) Generality. For different scenarios and complicated tasks, we might need to be redesigned the model policy manually for each new scenario.
To solve those problems, existing works such as Imitation Learning (IL) is most popular approach, which can learn a driving policy by collecting the expert driving data(Bojarski et al., 2016), (Codevilla et al., 2018), (Bansal et al., 2018) and (Chen et al., 2019a). However, those methods can suffer from the following shortcomings for imitation learning: (1) High training cost and sample complexity. (2) Conservation of driving policy. Due to the collect driving data from the human expert, which can only learn driving skills that are demonstrated in the datasets. (3) Limitation of driving performance. The performance of imitation learning policy is impossible exceed human-level performance.
What’s more, the driving policy based on Reinforcement Learning (RL) is also popular method in recent years, which can automatically learn and explore without any human expert data in various kinds of different driving cases, and it is possible to have a better performance than imitation learning. (Lillicrap et al., 2015) introduced a continuous control deep RL algorithm that trains a deep neural network policy for autonomous driving on a simulated racing track. (Wolf et al., 2017) used Deep Q-Network to learn to steer an autonomous vehicle to keep in the track in simulation. (Chen et al., 2018) developed a hierarchical deep RL framework to handle driving scenarios with intricate decision-making processes, such as navigating traffic lights. (Kendall et al., 2019) marked the first application of deep RL in real-world autonomous driving, where they trained a deep lane-keeping policy using only a single front-view camera image as input. (Chen et al., 2021) proposed an interpretable deep RL method for end-to-end autonomous driving. (Nehme & Deo, 2023) proposed safe navigation.(Murdoch et al., 2023) propose a partial end-to-end algorithm that decouples the planning and control tasks. (Zhou et al., 2023) proposes a method to identify and protect unreliable decisions of a DRL driving policy. (Zhang et al., 2023) a framework of constrained multi-agent reinforcement learning with a parallel safety shield for CAVs in challenging driving scenarios. (Liu et al., 2022) propose the Scene-Rep Transformer to enhance RL decision-making capabilities. (Huang et al., 2023) introduces a predictive behavior planning framework that learns to predict and evaluate from human driving data. In addition, combination of RL with Demonstrations are also learned more complex tasks such as dexterous manipulator. The techniques outlined in (Theodorou et al., 2010), (Van Hoof et al., 2015) and (Rajeswaran et al., 2017) have proven to be efficient in merging demonstrations and RL for improving learning speed. (Huang et al., 2024) propose an enhanced human in-the-loop reinforcement learning method, while they rely on human expert performance and can only accomplish simple scenario tasks. However, the existing methods also have some weakness: (1) They can only deal with the simple driving scenario tasks. (2)High costs of learning and exploration. (4) The safety and generalization of intelligent vehicles need further improvement. (4) Such these RL frameworks are mostly based on Partially Observable Markov Decision Process(POMDP)(Åström, 1965),(Hollinger & Sukhatme, 2014),(Agha-Mohammadi et al., 2014),(Indelman et al., 2015),(Karl et al., 2016). However, they cannot accurately predict future driving trajectories, and the accurate prediction of future driving trajectories is a precondition for making optimal decisions.
Figure 2: (a) RL Agent learns potential dynamics from past experience datasets of autonomous vehicle. (b) RL Agent predicts driving behavior in an imaginary space. (c) RL Agent interacts with driving environment.
In conclusion, the existing IL and RL methods can not well solve the safety, sampling efficiency and generalization of autonomous vehicles. To solve above problems, in this paper, we introduce a Safe and Generalized end-to-end Autonomous Driving System (SGADS) for complex and various scenarios. Our SGADS incorporates variational inference with normalizing flows, enabling the intelligent vehicle to accurately predict future driving trajectories. Moreover, we design safety constraint for intelligent vehicle. We further propose a behavior cloning as to augment search process of agent. We also contribute a demonstration dataset by having human experts drive using the G29 steering wheel. The experimental result shows our SGADS greatly augments safety performance, exhibits strong generalization in multiple testing maps, and enhances the training efficiency of intelligent vehicles in intricate urban scenarios compared to the existing methods. Moreover, we designed and explored up to 12 different data input types to evalutate their impact on the safety performance of intelligent vehicles. To sum up, our contributions are:
  • We introduces a novel SGADS designed to address the limitations of existing methods in solving safety, sampling efficiency, generalization problems in autonomous driving system.
  • SGADS incorporates variational inference with normalizing flows, enabling the intelligent vehicle to accurately predict future driving trajectories.
  • The experimental result shows that SGADS is capable of greatly enhancing safety, generalization, and training efficiency for intelligent vehicles.
  • We provide a dataset collected from human expert using the G29 steering wheel. Moreover, we have designed and explored 12 inputs to evalutate their influence on the safety performance of intelligent vehicles.
2 Problem formulation
2.1 System Framework
The proposed framework of our SGADS in Figure 1. Firstly, human expert collects the demonstrations offline using the Logitech G29 steering wheel. Then, the expert demonstrations are used as the initial experience replay buffers for training the entire model. Subsequently, a pre-training process is conducted to establish a model with human expert experience that do not update environmental data during training. The resulting model with human expert experience is then utilize to initialize network parameters for RL policy. Additionally, safety constraint is designed for intelligent vehicle, enhancing their safety performance. In addition, we design and explore 12 different types of images as the model inputs as shown in Figure 3 and Section 4.7.
2.2 Problem Formulation
Visual control(Thrun, 1999),(Silver & Veness, 2010), (Bengtsson et al., 2008) can be defined as a POMDP, the goal of the RL agent is to maximize expected cumulative reward $E[\sum_{t=0}^{\infty}\gamma_{t}r_{t}]$,$\gamma$ is the discount factor. Inspired by (Hafner et al., 2019a), (Hafner et al., 2019b), the RL agent in our SGADS enable the intelligent vehicle to predict future driving trajectories, which mainly consists of the following three parts:
(1) RL Agent learns potential dynamics from past experience datasets of autonomous vehicle to predict future rewards from actions and past observations. As shown in Figure 2(a), the entire learning process is a POMDP decision chain, using $p$ to represent prior probability, $q$ to represent posterior probability, agent learns to encode observations $o$ and actions $a$ into compact latent states $s$, and $\hat{o}_{t}$ is reconstructed with probability $q(\hat{o}_{t}|s_{t})$ while $s_{t}$ is determined via probability $p(s_{t}|s_{t-1},a_{t-1},o_{t})$.
(2) RL Agent predicts state values and actions of driving behaviors in an imaginary space, which maximize future value predictions by propagating gradients back through imagined trajectories. As shown in Figure 2(b), RL Agent is in a close latent state space where it can predict state $\hat{v}_{t}$, reward $\hat{r}_{t}$ and action $\hat{a}_{t}$ based on current input $o_{t-1}$ with probabilities $q(\hat{v}_{t},\hat{r}_{t},\hat{a}_{t}|s_{t})$, $p(s_{t}|s_{t-1},\hat{a}_{t-1})$, $q(\hat{a}_{t-1}|s_{t-1})$.
(3) RL Agent encodes the history of the episode to compute the current model state and predict the next action to execute in the environment. As shown in Figure 2(c), RL Agent predicts next action values $\hat{a}_{t+1}$ by encoding historical trajectory information via calculating probabilities $q(\hat{a}_{t+1}|s_{t+1})$, $p(s_{t+1}|s_{t},a_{t},o_{t+1})$.
Figure 3: Multiple types of input images
3 Methodology
3.1 Model Learning
We introduces a novel SGADS model consists of components can be constructed the PGM of POMDP as follow:
State transition model: $\quad p_{\theta}(s_{t}|s_{t-1},a_{t-1})$ (1)
Reward model: $\quad p_{\theta}(r_{t}|s_{t})$
Observation model: $\quad p_{\theta}(o_{t}|s_{t})$
Belief inference model: $\quad q_{\psi}(s_{t}|\tau_{t},o_{t})$
Where $p$ is prior distribution, $q$ is posterior distribution, $\tau_{t}=\{o_{1},a_{1},\cdots,o_{t-1},a_{t-1},o_{t},a_{t}\}$. The above of components can be optimized by maximizing the Evidence Lower BOund (ELBO)(Jordan et al., 1998),(De Cao et al., 2020):
$\mathcal{M}_{\text{model}}(\theta,\psi,\omega)=\sum_{t=1}^{T}\big{(}\mathbb{E}_{q(s_{t}\mid o_{\leq t},a_{<t})}[\log p(o_{t}\mid s_{t})+$ (2)
$\log p(r_{t}\mid s_{t})]-\big{.}{\mathbb{E}_{q(s_{t-1}\mid\tau_{t-1},o_{t-1})}}[D_{\rm KL}\big{(}q(s_{t}\mid\tau_{t},o_{t})\|$
$p(s_{t}\mid s_{t-1},a_{t-1})]\big{)}\big{)}$
In the continuous and dynamic space, existing methods based on Normalizing Flows(NF)(Dinh et al., 2016) , (Huang et al., 2018), (Rezende & Mohamed, 2015) can learn more flexible and generalized beliefs. These methods provide a solid foundation for RL agents based on POMDP to accurately predict future driving trajectories. Inspired by (Chen et al., 2022), we can substitute the probability density with Normalizing Flow in the KL-divergence term of Equation 2.
$q_{K}(s_{t}|\tau_{t},o_{t})=\log q_{0}(s_{t}|\tau_{t},o_{t})-\sum_{k=1}^{K}|det\frac{\partial f_{\theta_{k}}}{\partial s_{t,k-1}}|$ (3)
$p_{K}(s_{t}|\tau_{t})=\log p_{0}(s_{t}|\tau_{t})-\sum_{k=1}^{K}|det\frac{\partial f_{\omega_{k}}}{\partial s_{t,k-1}}|$
Where $q_{K}=q_{\psi,\theta}$, $p_{K}=q_{\psi,\omega}$. The final inferred belief is obtained by propagating $q_{\psi}(s_{t}|\tau_{t},o_{t})$ through a set of NF mappings denoted $f_{\theta_{K}}\dots f_{\theta_{1}}$ to get a posterior distribution $q_{\psi,\theta}(s_{t}|\tau_{t},o_{t})$. The final prior is obtained by propagating $p_{\psi}(s_{t}|\tau_{t})$ through a set of NF mappings denoted $f_{\omega_{K}}\dots f_{\omega_{1}}$ to get a prior distribution $p_{\psi,\omega}(s_{t}|\tau_{t})$.
3.2 Policy Learning
If the initialization of the policy is poor, the learning process may be very slow and the algorithm will explore states and actions that are irrelevant to the task. To address this problem, we use behavior cloning to provide a policy $\pi^{*}$ via expert demonstrations and then to train a model $\mathcal{M}_{\text{expert}}$ with some expert ability. The loss function for training the model in the first stage is as follows:
[…]
In the second stage of the model, we employ $\mathcal{M}_{\text{expert}}$ to initialize a model trained by deep reinforcement learning policies, which greatly reduces the sampling complexity of the deep RL policy intelligent vehicle. The training loss of the actor model in the second stage is as follows:
${\rm\mathcal{M}_{actor}(\phi,\eta)}=\mathbb{E}\Big{(}\sum_{i=1}^{N}\sum_{\tau=t}^{t+H}\left(V_{i,\tau}^{\lambda}\right)\Big{)}+k\ln\pi^{*}_{\eta}(a_{t}^{\prime}|s_{t}^{\prime})$ (5)
Where $k$ represents the balance between the cloning policy and deep reinforcement learning policy, and is set as a constant based on empirical data. Then we can train the critic to regress the TD($\lambda$) (Sutton & Barto, 2018) target return via a mean squared error loss:
$\mathcal{M}_{\text{critic }}(\xi)=\mathbb{E}\Big{[}\sum_{i=1}^{N}\sum_{\tau=t}^{t+H}\frac{1}{2}\left(v_{\xi}\left(s_{i,\tau}\right)-V_{i,\tau}^{\lambda}\right)^{2}\Big{]}$ (6)
The $v_{\xi}\left(s_{\tau}\right)$ is designed to forecast the cumulative discounted sum of future rewards achievable by the actor, starting from an initial state $s_{\tau}$. $V_{\tau}^{\lambda}$ can be defined as follow:
$V_{\tau}^{\lambda}=(1-\tau)v_{\xi}(s_{\tau+1})+\lambda V_{\tau+1}^{\lambda},\quad\tau\textless{t+H}$ (7)
Where $\xi$ denote the parameters of the critic network and $H$ is the prediction horizon. Finally, the loss function for the second stage training of the whole model is as follows:
$\underset{\psi,\xi,\phi,\theta,\omega,\eta}{\rm min}\alpha_{0}{\rm\mathcal{M}_{critic}(\xi)}-\alpha_{1}{\rm\mathcal{M}_{actor}(\phi,\eta)}-\alpha_{2}{\rm\mathcal{M}_{model}(\theta,\psi,\omega)}$ (8)
We jointly optimize the parameters of model loss $\psi,\theta,\omega$, critic loss $\xi$ and actor loss $\phi,\eta$, where $\alpha_{0},\alpha_{1},\alpha_{2}$ are coefficients for different components.
3.3 Safety constraint
Figure 4: Randomly sample sensor inputs Lidar_noground $o_{1},o_{2},\cdots,o_{15}$, and then our model can imagine driving behaviors $\hat{o}_{16},\hat{o}_{17},\cdots,\hat{o}_{30}$. The results show that our model can make precise and diverse predictions with no mode mixup and less blur.
To keep the intelligent vehicle running safely and smoothly in complex environments, we add safety and robustness constraints to the reward function of the algorithm(Chen et al., 2019b), including time to collision and smooth steering.
(1)Front time to collision. When around vehicles are within the distance of ego vehicle (our agent vehicle) head in our setting, then we can calculate the front time to collision between ego vehicle and around vehicles. Firstly, the speed and steering vector $(s_{\tau},a_{\tau})\in\mathbb{S}$ of the ego vehicle are defined, where $s_{\tau}$ represents the angle vector of vehicle steering and $a_{\tau}$ represents the acceleration vector of the vehicle in local coordinate system. Secondly, two waypoints closest to the current ego vehicle are selected from the given navigation routing as direction vectors $\vec{w_{p}}$ for the entire route progression, where $\rightarrow$ indicates a vector in world coordinates. The position vectors for both ego vehicle and around vehicles are represented by $(\vec{x_{t}}^{*},\vec{y_{t}}^{*})$, respectively. Finally, $\delta_{e}$ and $\delta_{a}$ representing angles between position vectors for ego vehicle and around vehicles with respect to $\vec{w_{p}}$ are calculated respectively.
$\vec{w_{p}}=\big{[}(\frac{\vec{w_{t+1}^{x}}-\vec{w_{t}^{x}}}{2})-(\vec{w_{t}^{x}}),(\frac{\vec{w_{t+1}^{y}}-\vec{w_{t}^{y}}}{2})-(\vec{w_{t}^{y}})\big{]}$ (9)
$\delta_{e}=\frac{[\vec{v_{t}^{x*}},\vec{v_{t}^{y*}}]\cdot\vec{w_{p}}}{\|\vec{v_{t}^{x*}},\vec{v_{t}^{y*}}\|_{2}\quad\|\vec{w_{p}}\|_{2}},\delta_{a}=\frac{[\vec{v_{t}^{x}},\vec{v_{t}^{y}}]\cdot\vec{w_{p}}}{\|\vec{v_{t}^{x}},\vec{v_{t}^{y}}\|_{2}\quad\|\vec{w_{p}}\|_{2}}$
Where, $l$ is the length of the set of waypoints $\mathbb{W}$ stored. The variable $t\in\tau$, and $\vec{w_{t}^{x}}\in\mathbb{W}_{1}$ represents the $x$ coordinate of the first navigation point closest to the intelligent vehicle on its current route at time $t$. Similarly, $\vec{w_{t+1}^{x}}\in\mathbb{W}_{2}$. Furthermore, it is possible to calculate the $F_{ttc}$ as follows:
$F_{ttc}=\frac{\|\vec{x_{t}}-\vec{x_{t}^{*}},\vec{y_{t}}-\vec{y_{t}^{*}}\|_{2}}{\left|\|\vec{v_{t}^{x*}},\vec{v_{t}^{y*}}\|_{2}sin(\delta_{e})-\|\vec{v_{t}^{x}},\vec{v_{t}^{y}}\|_{2}sin(\delta_{a})\right|}$ (10)
(2)Lateral time to collision. When around vehicles are not within the distance of ego vehicle head in our setting, we consider significantly the $L_{ttc}$. The calculation method for $L_{ttc}$ and $F_{ttc}$ is the same. However, the collision constraint effect of $L_{ttc}$ on intelligent vehicle is limited, mainly due to the slow reaction time of intelligent vehicle to $L_{ttc}$, lack of robustness and generalization ability. Therefore, we have implemented a method of assigning values to different intervals for $L_{ttc}$ as follows:
$&\text{min}(z_{\tau},c_{\tau}+1.0),\\ &\nu_{g}\leq(c_{lttc}-1.5)\text{ and }\mu_{a}\leq(c_{\tau}-0.5).\\ &\text{min}(z_{\tau},c_{\tau}-1.8),\\ &\nu_{g}\leq(c_{\tau}-3.0)\text{ and }\mu_{a}\leq(c_{\tau}-2.0).\\ &\text{min}(z_{\tau},c_{\tau}-3.0),\\ &\nu_{g}\leq(c_{\tau}-3.5)\text{ and }\mu_{a}\leq(c_{\tau}-3.0).\\ \end{aligned}\right.\vspace{-0.2cm}$ (11)
Where $c_{\tau}$ is the empirical const of $L_{ttc}$ in our setting at (5,7), $z_{\tau}$ is the ttc based on their combined speed. $\nu_{g}$ is the ttc obtained by calculating the longitudinal velocity. $\mu_{a}$ is the ttc obtained by calculating the lateral velocity.
(3) Smooth steering is defined as $|s_{t}^{\delta}-s_{t}^{*\delta}|\in e_{c}$. $s_{t}^{\delta}$ is the actual steering angle. $s_{t}^{*\delta}$ is the predicted steering angle based on policy $\pi$. The range of $e_{c}$ can be established based on empirical data.
4 Experiment
Figure 5: The road networks of the CARLA include routes for Town01, Town02, Town03, Town04, Town05, and Town06, as well as more complex scenarios in Town03, Town04, and Town05
The model training in this paper is deployed on systems including Python=3.6, Pytorch-gpu=1.5.0, GeForce RTX-3090 and CARLA=0.9.10 running on Linux. In order to evaluate the performance of our SGADS in addressing challenges related to sampling efficiency, safety, generalization and prediction for future driving trajectories, we respectively designed corresponding experiments to answer each of these questions. Additionally, we formulated experiments involving 12 different sets of input data to evaluate the influence on the safety of autonomous vehicles.
Table 1: In the training stage, the steps and times for all methods are compared under different baseline average collision distances in Town03(max episodes length=500), the number of vehicles is 200, and calculating Average Driving Distance(Avg Dis) with 10 episodes. Particularly, +$\infty$+ ∞ means the situation where the training time has exceeded 250hours and the model has still not reached the baseline.
Method Input Avg dis=50m Avg dis=150m Avg dis=200m
$\downarrow$ $\downarrow$ $\downarrow$ times $\downarrow$ $\downarrow$ times $\downarrow$
DDPG Lidar +$\infty$+ ∞ +$\infty$+ ∞ +$\infty$+ ∞ +$\infty$+ ∞ +$\infty$+ ∞ +$\infty$+ ∞
SAC Lidar +$\infty$+ ∞ +$\infty$+ ∞ +$\infty$+ ∞ +$\infty$+ ∞ +$\infty$+ ∞ +$\infty$+ ∞
TD3 Lidar $\geq$161 $\geq$192h +$\infty$+ ∞ +$\infty$+ ∞ +$\infty$+ ∞ +$\infty$+ ∞
DQN Lidar $\geq$163 $\geq$53h +$\infty$+ ∞ +$\infty$+ ∞ +$\infty$+ ∞ +$\infty$+ ∞
Latent_SAC Lidar $\geq$167 $\geq$40h $\geq$300 $\geq$212h +$\infty$+ ∞ +$\infty$+ ∞
SGADS Lidar $\geq$21 $\geq$1.3h $\geq$121 $\geq$20h $\geq$321 $\geq$48h
SGADS Lidar_ng $\geq$15 $\geq$0.9h $\geq$139 $\geq$9h $\geq$211 $\geq$38h
4.1 Training and Evaluation Setting
4.1.1 Traning environment
The training and evaluation of our entire autonomous driving perception-decision-making joint system are conducted on the CARLA(Dosovitskiy et al., 2017). In order to comprehensively evaluate the performance of our SGADS, we utilized six maps in CARLA ranging from Town01 to Town06, as illustrated in Figure 5. Town01, a small, simple town with a river and several bridges. Town02, a small simple town with a mixture of residential and commercial buildings. Town04, a small town embedded in the mountains with a special infinite highway. Town05, squared-grid town with cross junctions and a bridge. Town06, long many lane highways with many highway entrances and exits. Particularly, Town03 is the most complex town with a 5-lane junction, a roundabout, unevenness, a tunnel, and more. To evaluate more realistically, we adopted the approach of randomly generating driving routes, resetting the coordinates of all vehicles in the map, and randomizing the driving scenarios in each episode.
Figure 6: (a) CARLA connects with the G29 steering wheel (b) Human expert collects the datasets via the G29 steering wheel
4.1.2 Measure Safety
Safety is defined as the conditions in which an autonomous vehicle in its environment can follow a predetermined route without colliding with surrounding vehicles, avoid major deviations from a predetermined trajectory, avoid stopping on the road surface, avoid contact with a roadside wall or guardrail, and successfully navigate a complex intersection or roundabout without a collision. In this experimental environment setting, if the intelligent vehicle fails to comply with the above regulatory constraints, it will be penalized, resulting in the termination of the current episode. In addition, when an episode ends, all scenes and vehicles are forcibly reset, and the reset of the intelligent car is also considered as the vehicle experienced a collision, resulting in the end of the episode. Therefore, we measure the safety using the Average Driving Distance(Avg Dis), which you can think of as the average distance a self-driving car can safely travel in multiple episodes. When intelligent vehicle travels a longer driving distance within the episode, it indicates that its safety performance is higher. Simultaneously, as the episode length increases and our testing extends, the vehicle is more likely to have a longer safe driving distance. Additionally, we evaluate the ability to safely travel the maximum distance over multiple episodes using Max Driving Distance(Max Dis), calculated as the maximum value of the driving distance across multiple episodes.
4.2 Human expert collects teaching datasets
CARLA can be operated and controlled through using the python API. Figure 6 shows that we establish a connection between the Logitech G29 steering wheel and the CARLA, and then human expert can collect the datasets of teaching via the G29 steering wheel. Specifically, we linearly map accelerator pedals, brake pedals, and turning angles into accel[0,3](min,max), brake[-8,0] (min,max), steer[-1,1](left,right). The tensors are written into user-built Python scripts and combined with CARLA built-in Python API so that users can provide input from their steering wheels to autonomous driving cars in CARLA simulator for $\mathcal{D}_{expert}$ collection. Particularly, we contributed a dataset collected through human expert steering wheel control.
4.3 Evaluate the prediction of driving trajectories
In order to accurately evaluate our model prediction of driving behavior for intelligent vehicle in the imagination space, this problem can be viewed as a special POMDP problem with the reward value maintained at 0. Then, two scenarios of original sensor inputs for intelligent vehicle at time steps 1-15 are randomly sampled as shown in Figure 4, followed by further allowing our model to predict the driving behaviors of intelligent vehicle at time steps 16-30 under an imagination space length of $H=15$. The result shows that our model can accurately predict future driving trajectories.
Figure 7: All training for SGADS was completed in Town03, followed by separate evaluations from Town01 to Town06. In all evaluations, the calculations for Avg Dis and Max Dis are based on 5 episodes. Due to the smaller size of the map in town02, the number of vehicles is limited to 50, while in other maps, the number of vehicles is set to 200.
4.4 Evaluate Sampling efficiency
In order to evaluate the performance of our autonomous driving system more effectively, we have conducted various comparisons with existing methods such as DDPG(Lillicrap et al., 2015), SAC(Haarnoja et al., 2018), TD3(Fujimoto et al., 2018), DQN(Mnih et al., 2015) and Latent_SAC(Chen et al., 2021). These methods achieve their best performance when using lidar as input: As shown in Table 1, compared to other methods, our SGADS achieves at least a 30 times and 10 times efficiency improvement in training time when reaching baselines of 50 meters and 150 meters, respectively, with Lidar as input. Particularly, when the baseline is set at 200 meters and the training time exceeds 250 hours, only our model achieves this baseline. Additionally, by exploring the input as Lidar_ng, efficiency of the model is further enhanced compared to Lidar. This indicates that our SGADS can effectively reduce training costs and improve sampling efficiency compared to existing methods.
4.5 Evaluate Safety
To comprehensively evaluate safety, we chose the most complex urban map, Town03. Our SGADS is initially trained in Town03 and subsequently evaluated in Town03 under random and complex scenarios. In random scenarios, the ego vehicle and around vehicles can appear at any location on the map. In fixed scenarios, they are restricted to random appearances within a specific scene range. The difference between the two lies in the range of random selection, but both follow the randomization mechanisms of CARLA in every episode during training and evaluation.
As shown in Figure 7, the first row from left to right illustrates the three scenarios we evaluated in Town03, including random driving scenarios, fixed complex roundabouts, and fixed intersections. The experimental results show that SGADS exhibits outstanding Avg Dis and Max Dis in both random driving scenarios and complex scenarios, significantly surpassing existing methods. This further emphasizes that our approach has the potential to significantly enhance the safety of intelligent vehicles.
Additionally, to clarify the sources of safety benefits, we decomposed SGADS into three components, variational inference based on normalizing flows (Base), safety constraints (SC), and behavior cloning (BC). We then conducted evaluations using three comparison settings, Base, Base+SC, and Base+SC+BC. This is illustrated in the first row of evaluations in Figure 7 for the ablation evaluation in Town03. Under the input conditions of lidar or Lidar_noground, the safety benefits of SGADS primarily result from the design of the SC component, while the BC component can further provide additional benefits.
4.6 Evaluate Generalization
To evaluate the generalization performance of SGADS, in each episode of the experiment, scenes are randomly set, driving routes are generated randomly, and environmental vehicles are randomly generated, ensuring the complexity and diversity of the evaluation. As shown in the second row of Figure 7, we initially train SGADS and other methods in Town03 and evaluate respectively in Town01, Town02, Town04, Town05, and Town06. The experimental results demonstrate that SGADS exhibits significantly superior Avg Dis and Max Dis compared to other methods, and the required training time is relatively short, confirming its efficient training performance.
Additionally, despite the relatively small size of the Town02 map, resulting in a smaller upper limit for driving distance, SGADS still outperforms other methods in terms of Avg Dis, Max Dis, and training efficiency. This indicates that SGADS exhibits outstanding generalization capabilities, primarily benefiting from the integration of various aspects, including realistic random mechanisms close to the real world in CARLA, trajectory prediction based on normalizing flows, safety constraints, and the fusion of human expert demonstrations.
Table 2: Evaluate the performance of intelligent vehicle under multiple input images in town03, training steps=100k. Avg Dis is based on 5 episodes, the number of vehicles is 200.
Multiple input Avg dis(m)
birdeye 29.4
lidar 38.6
camera 41.5
lidar+camera 56.4
semantic 26.3
depth 29.4
lidar_ng 64.7
multi-fusion1 52.1
multi-fusion2 47.2
lidar+depth _ng 32.5
lidar_ng+multi-fusion3 36.4
lidar_ng +camera_gray 48.8
4.7 Multiple types of input images
We compare various input image types for evaluating the performance of intelligent vehicle based on BC of the SGADS, as shown in Table 2. The representation of multiple inputs is described in the section Appendix A.1 Avg Dis of lidar_noground reaches the highest value compared with all other input types. This is because lidar_noground removes a large amount of redundant information, reduces the difficulty of world model understanding environment semantics, and also involves stationary status of intelligent vehicle in experiment. The results show that the lidar_noground input is relatively optimal. However, it is worth noting that the effects of these 12 different input types are relatively small, with the Avg Dis e only varying between 20 and 40 meters. This shows that different data types have a minimal impact on the safety performance of intelligent vehicles.
5 Conclusion
In order to address the challenges of low sample efficiency, low safety, limited generalization, and accurate prediction of future driving trajectories in autonomous driving systems, we propose a safe, efficient, and generalizable end-to-end autonomous driving system (SGADS). Experimental results demonstrate that compared to existing methods, the safety constraints in SGADS significantly enhance the safety of intelligent vehicles, and demonstrations further improve safety and sample efficiency. Variational inference based on normalizing flow accurately predicts future driving trajectories, providing a precondition for optimal decision making. Finally, the fusion of these three modules greatly enhances the generalization capabilities of intelligent vehicles, significantly surpassing existing methods. Additionally, we observe that despite differences in input data types, these variances have minimal impact on the safety. We contribute a demonstration dataset by having human experts drive using the G29 steering wheel. However, there is still space for improvement in aligning our SGADS system with real-world safety and generalization requirements.
Appendix A Appendix
A.1 Multiple types of input images
The 12 types of input data we designed are mainly categorized into single-modal and single-image input, single-image and multimodal fusion, and multiple images and multimodal fusion.
1)Single-modal and input of a single image. As Figure 3 shown, the lidar images, which project the 3D point cloud information from lidar onto a 2D point cloud image, with each pixel color determined by whether there is lidar or other relevant pixel information on the corresponding area. Navigation path is rendered in blue and surrounding road conditions are represented by green rectangular boxes to indicate participants such as vehicles, pedestrians etc. Particularly, lidar_noground is created to remove redundant ground truth information from the 2D point cloud image. Moreover, we also consider camera, semantic, birdeye and depth as our sensor inputs.
2)Single-image and multimodal fusion. The input of single-image and multi-modal fusion involve fusing lidar, rgb forward-facing grayscale image(camera_gray), and navigation path into a composite rgb image with three types of information. The fused image has three channels, multi-fusion1(lidar,camera_gray,routing). Similarly having multi-fusion2(lidar,depth,routing) and multi-fusion3(lidar, depth,0).
3)Multiple images and multi-modal fusion. Multiple fusion can complement the shortcomings of a single input source and provide richer and more effective information. Therefore, we also design several single-modal fusion inputs as shown on the right side of Figure 3, including lidar_noground and multi_fusion3, lidar_noground and depth, lidar-noground and camera_gray, as well as camera and lidar.
A.2 Reward function
We use the following reward function in our experiments:
$\text{R}=200\cdot r_{c}+200\cdot r_{ft}+50\cdot r_{lt}+2\cdot r_{sc}+v_{lon}+10\cdot r_{f}+r_{o}-5\alpha^{2}+0.2\cdot r_{lat}-0.1$ (12)
where $r_{c}$ is the reward related to collision, which is set to -1 if the ego vehicle collides and 00 otherwise. $v_{lon}$ is the longitudinal speed of the ego vehicle. $r_{f}$ is the reward related to running too fast, which is set to $-1$ if it exceeds the desired speed (8 m/s here) and 00 otherwise. $r_{o}$ is set to $-1$ if the ego vehicle runs out of the lane, and 00 otherwise. $\alpha$ is the steering angle of the ego vehicle in radians. $r_{lat}$ is the reward related to lateral acceleration, which is calculated by $r_{lat}=-|\alpha|\cdot v_{lon}^{2}$. The last constant term is added to prevent the ego vehicle from standing still. $r_{ft}$ represents the time to collision in the forward direction. $r_{lt}$ represents the time to collision in the lateral direction. $r_{sc}$ represents the smoothness constraint.
A.3 Hyperparameter settings
$\mathcal{M}_{model}$, the KL regularizer is clipped below 3.0 free nats for imagination range $H=15$ using the same trajectories for updating action and value models separately with $\lambda=0.99$ and $\lambda=0.95$, while $k=1.5$. The size of all our trainig and evaluating images is $128\times 128\times 3$. For Algorithm 1, a random seed $S=5$ is used to collect datasets for the $ego$ vehicle before updating the model every $C=100$ steps during training process.
Table 3: Hyperparameter settings for the training and evaluation of each baseline
Method batch size model size eval episodes action repeat
DDPG 256 32 5 2
SAC 256 32 5 2
TD3 256 32 5 2
DQN 256 32 5 2
Latent_SAC 256 32 5 2
SGADS 32 32 10 1
Table 4: Hyperparameter settings for the learning rate of each baseline
Method model learning rate actor learning rate value learning rate
DDPG $1\times 10^{-4}$ $3\times 10^{-4}$ $3\times 10^{-4}$
SAC $1\times 10^{-4}$ $3\times 10^{-4}$ $3\times 10^{-4}$
TD3 $1\times 10^{-4}$ $3\times 10^{-4}$ $3\times 10^{-4}$
DQN $1\times 10^{-4}$ $3\times 10^{-4}$ $3\times 10^{-4}$
Latent_SAC $1\times 10^{-4}$ $3\times 10^{-4}$ $3\times 10^{-4}$
SGADS $1\times 10^{-3}$ $8\times 10^{-5}$ $8\times 10^{-5}$
A.4 The world model reconstructs the input images from the original sensors
We explores the differences between input images from original sensors and the corresponding reconstructed input images from a world model for 8 types of input. As shown in Figure 8, multiple comparisons are made between the reconstructed input types generated by the world model and their corresponding original sensor inputs. Among them, multi-fusion2, lidar_noground, lidar+camera and lidar reconstructions are very clear and highly consistent, indicating that $q(o_{t}|s_{t})$ has a precise decoding capability without causing loss of $s_{t}$. However, birdeye, semantic, (lidar_noground and multi-fusion3), and (lidar_noground and camera_gray) of reconstructions are not as clear as their sensor input. This suggests that world model have difficulty understanding large amounts of irrelevant information related to driving tasks resulting in unclear reconstruction outputs.
Figure 8: Randomly sampled frames to reconstruct the input images from the original sensors of SGADS on 8 types of input. For each type of image, first row: original sensor inputs. Second row: reconstructed images.
A.5 Algorithm
we propose the framework of safe and generalized end-to-end autonomous driving system with reinforcement learning and demonstrations is presented in Algorithm 1.
Algorithm 1 SGADS
Input: batch size $B$, batch length $L$, imagination horizon $H$, step $T$, trajectory $\tau$, number of trajectories $N$, initialize datasets $\mathcal{D}$ with seeds.
1:  The first stage of trainingCollect $\mathcal{D}_{expert}$ via the G29 steering.Randomly initialize parameters $\theta,\phi,\psi,\eta,\omega$.
2:  while $\mathcal{M}_{expert}$ does not converge do
3:     for $c=1,2,\ldots,C$ do
4:        Randomly sample $B$ sequences $\{(a_{t}^{\prime},o_{t}^{\prime},r_{t}^{\prime})\}_{t=k}^{k+L}$ from $\mathcal{D}_{expert}$.
5:        Calculating the states $p_{\eta}(s_{t}^{\prime}|s_{t-1}^{\prime},a_{t-1}^{\prime},o_{t}^{\prime})$.
6:        Update parameters $\phi,\eta$.
7:     end for
8:  end while
9:  The second stage of trainingInitialize all neural network parameters using $\mathcal{M}_{expert}$.
10:  while $\mathcal{M}_{model}$ does not converge do
11:     for $c=1,2,\ldots,C$ do
12:        Ramdomly sample $B$ sequences from $\mathcal{D}$.
13:        Infer belief state $s_{t}\sim q_{K}(s_{t}|s_{t-1},a_{t-1},o_{t})$.
14:        for $i=1,2,\ldots,N$ do
15:           Rollout imaginary trajectories ${(s_{i,\tau},a_{i,\tau})}_{\tau=t}^{t+H}$ with belief transition model.
16:        end for
17:        Update $\theta,\phi,\psi,\xi,\eta,\omega$.
18:     end for
19:     for $t=1,2,\ldots,T$ do
20:        Compute $s_{t}\sim q_{K}(s_{t}|s_{t-1},a_{t-1},o_{t})$ from history.
21:        Compute $a_{t}\sim\pi(a_{t}|s_{t})$ with action model.
22:        $r_{t},o_{t+1}\leftarrow$ TTC and CSS.
23:     end for
24:     Update $\mathcal{D}\leftarrow\mathcal{D}\cup\{(a_{t},o_{t+1},r_{t})_{t=1}^{T}\}$
25:  end while
A.6 Further comparisons were made regarding safety and generalization
In the existing baseline methods, the default number of environment vehicles is set to 100. We also conducted comparisons in terms of generalization and safety as shown in Figure 9 and 10. The experiments indicate that our model significantly outperforms existing methods.
Figure 9: The model is trained in Town03 and then evaluated for generalization in Town01, Town04 and Town05.
Figure 10: The model is trained in Town03 and then evaluated for safety in roundabout of the Town03, evluated for generalization in Town06.
A.7 More results regarding predictions of future driving trajectories
The accurate prediction of future driving trajectories is a precondition for making optimal decision making. Random samples of driving trajectories for the first 15 time steps were collected from the sensor. Subsequently, the model predicted the driving trajectories for the next 15 time steps, and the ground truths for these trajectories were also provided We provide additional results regarding predictions of future driving trajectories as shown in Figure 10 …Figure 18.
(a) Randomly sample ground truth of inputs Lidar $o_{1},o_{2},\cdots,o_{15}$
(b) Randomly sample ground truth of inputs Lidar $o_{16},o_{17},\cdots,o_{30}$
(c) Our model can imagine driving behaviors $\hat{o}_{16},\hat{o}_{17},\cdots,\hat{o}_{30}$
Figure 11: We randomly sampled input images, and then SGADS was used to make predictions
(a) Randomly sample ground truth of inputs Lidar $o_{1},o_{2},\cdots,o_{15}$
(b) Randomly sample ground truth of inputs Lidar $o_{16},o_{17},\cdots,o_{30}$
(c) Our model can imagine driving behaviors $\hat{o}_{16},\hat{o}_{17},\cdots,\hat{o}_{30}$
Figure 12: We randomly sampled input images, and then SGADS was used to make predictions
(a) Randomly sample ground truth of inputs Lidar $o_{1},o_{2},\cdots,o_{15}$
(b) Randomly sample ground truth of inputs Lidar $o_{16},o_{17},\cdots,o_{30}$
(c) Our model can imagine driving behaviors $\hat{o}_{16},\hat{o}_{17},\cdots,\hat{o}_{30}$
Figure 13: We randomly sampled input images, and then SGADS was used to make predictions
(a) Randomly sample ground truth of inputs Lidar $o_{1},o_{2},\cdots,o_{15}$
(b) Randomly sample ground truth of inputs Lidar $o_{16},o_{17},\cdots,o_{30}$
(c) Our model can imagine driving behaviors $\hat{o}_{16},\hat{o}_{17},\cdots,\hat{o}_{30}$
Figure 14: We randomly sampled input images, and then SGADS was used to make predictions
(a) Randomly sample ground truth of inputs Lidar $o_{1},o_{2},\cdots,o_{15}$
(b) Randomly sample ground truth of inputs Lidar $o_{16},o_{17},\cdots,o_{30}$
(c) Our model can imagine driving behaviors $\hat{o}_{16},\hat{o}_{17},\cdots,\hat{o}_{30}$
Figure 15: We randomly sampled input images, and then SGADS was used to make predictions.
(a) Randomly sample ground truth of inputs Lidar $o_{1},o_{2},\cdots,o_{15}$
(b) Randomly sample ground truth of inputs Lidar $o_{16},o_{17},\cdots,o_{30}$
(c) Our model can imagine driving behaviors $\hat{o}_{16},\hat{o}_{17},\cdots,\hat{o}_{30}$
Figure 16: We randomly sampled input images, and then SGADS was used to make predictions
(a) Randomly sample ground truth of inputs Lidar $o_{1},o_{2},\cdots,o_{15}$
(b) Randomly sample ground truth of inputs Lidar $o_{16},o_{17},\cdots,o_{30}$
(c) Our model can imagine driving behaviors $\hat{o}_{16},\hat{o}_{17},\cdots,\hat{o}_{30}$
Figure 17: We randomly sampled input images, and then SGADS was used to make predictions
(a) Randomly sample ground truth of inputs Lidar $o_{1},o_{2},\cdots,o_{15}$
(b) Randomly sample ground truth of inputs Lidar $o_{16},o_{17},\cdots,o_{30}$
(c) Our model can imagine driving behaviors $\hat{o}_{16},\hat{o}_{17},\cdots,\hat{o}_{30}$
Figure 18: We randomly sampled input images, and then SGADS was used to make predictions